Find the equivalent functional definition for each of the following and draw a graph of the function.


f(x)= x+|x|.

1
Expert's answer
2022-06-09T13:05:43-0400

"f(x)= x+|x|."


As "|x|=\{\begin{matrix}
 x & \text{if }x\geq0, \\
 -x & \text{if }x<0,
\end{matrix}"

we have:

"f(x)=\{\begin{matrix}
 x+x & \text{if }x\geq0, \\
 x+(-x) & \text{if }x<0,
\end{matrix}"


"f(x)=\{\begin{matrix}
 2x & \text{if }x\geq0, \\
 0 & \text{if }x<0,
\end{matrix}"